OpenStudy (anonymous):

what is the ordered pair of y=4x? 1,3 -1,-4 -4,-1 1,-4 please i need help ill give you a medal

OpenStudy (anonymous):

you just need to fill in the x & the y. and see what doesnt work. its -1,-4 because -4=4(-1)

OpenStudy (anonymous):

thanks ashley are you sure

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

y= -7x+2 do you know the ordered pair for this one

OpenStudy (austinl):

This one is handy because you are given choices. The second one you just posted is a little more difficult if there are no options.

OpenStudy (anonymous):

is there options for this one too?

OpenStudy (anonymous):

yeah hold on

OpenStudy (anonymous):

1,2 8,-54 5,-35 1,-7

OpenStudy (austinl):

Plug them in and see which one makes the inequality true!

OpenStudy (anonymous):

its 8,-54

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

-54=-7(8)+2

OpenStudy (anonymous):

hold on if you can i got more hahaha

OpenStudy (anonymous):

okay!

OpenStudy (anonymous):

x-3=-5 -5/3 -2 -8 -15

OpenStudy (anonymous):

-2-3=-5 ( that one was the easiest...)

OpenStudy (anonymous):

so what the answer'

OpenStudy (anonymous):

the only one you had to fill in was the x. so its -2

OpenStudy (anonymous):

ok hold on there are like 25 questions and if i dont pass this test i literally go back to jail because my grade will drop down and when that happen im meat

OpenStudy (anonymous):

do you not understand it?

OpenStudy (anonymous):

no... ok 6.4 = 4x 2.4 1.6 4 6.4 im just glad i have people like you... very smart and intelligent

OpenStudy (anonymous):

its 1.6 all you have to do is fill in the numbers in the variable. so if you have a calculator.. 4 times 1.6 = 6.4

OpenStudy (anonymous):

thanks how do you give somebody a gold or whatever

OpenStudy (anonymous):

its just when you click best response

OpenStudy (anonymous):

oo

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

is there anything else you need help with?

OpenStudy (anonymous):

yes maam hold on just one second

OpenStudy (anonymous):

u there

OpenStudy (anonymous):

yep. just working on my own homework at the same time

OpenStudy (anonymous):

o

OpenStudy (anonymous):

do u have like a facebook this would be alot easier on there cause i could send you pics instead of typing it all out

OpenStudy (anonymous):

sure ill DM you the link

OpenStudy (anonymous):

ok